How many moles of N2 does a flask with a volume of 0.250 L at a pressure of 2.96 atm and a temperature of 250 K contain?
1
Expert's answer
2019-04-01T07:48:36-0400

"n(N2)=\dfrac{PV}{RT}=\dfrac{2.96atm*0.25L}{0.0821\dfrac{L*atm}{mol*K}*250K}=0.036 mol"